# Pattern formation in a coupled driven diffusive system

### v1

@@ II.3 Brief analysis of various regimes
The peak positions of the CF SF, which occur at $\left(q^{\ast},0\right)$, are listed in the second column, while the next three columns are associated with the $S_{-}$ values at $\left(q^{\ast},q_{y}\right)$, with $q_{y}\,L/2\pi=0,1$ and $2$. Note that the ratios of the entries in the third and fourth columns are dramatically different in the three cases: $\sim 1.4$, $38$, and $215$ for the microemulsion phase, the irregular stripes, and the regular stripes, respectively. Remarkably, once $q_{y}$ reaches $3\cdot 2\pi/L$, there is hardly any difference between the three rows! We also call attention to the $q^{\ast}$ of the irregular stripe regime, which is the same (for this $L$) as that of the disordered phase, but considerably smaller than that in the regular stripe regime, which may be also observed in Fig. 3 where the stripes at high density are visibly narrower. However, we still emphasize that, although $q^{*}$ takes the same value in both the disordered phase and the intermediate (irregular stripe) regime for $L=128$, it is independent of $L$ in the microemulsion phase. In Fig. 6, we plot the structure factors for $L=256$. The peak in the low density, disordered phase is well fit by a function of the form $A_{1}q^{2}\left[A_{2}q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$, consistent with the behavior observed in the DWRLG [1, 2]. Upon examination of other system sizes, e.g., $L=128$ and $192$, we consistently find $q^{*}=0.101\pm 0.001$ for the disordered regime.
Figure 6: Steady-state structure factor (SF) in the low-density regime using $L=256$ for $\rho=0.39$ and $\delta=1$, averaged over $25$ repetitions. Upper panels: (a) Charge fluctuation (CF) SF. (b) CF typical field configuration. Lower panels: (a) Density fluctuation structure factor (DF SF) (b) DF typical field configuration. The black circles (red diamonds) are the SF parallel (perpendicular) to the drive direction, $S_{\pm}(q_{x},0)$ and $S_{\pm}(0,q_{y})$, respectively. Error bars are smaller than the symbols. The blue line is the fit $A_{1}\,q^{2}[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}]^{-1}$ using 2020 points around the peak. Analogous behavior is found for $L=128,192$. We find $q^{*}=0.101(1)$ for $L=128,192$ and $256$. Observe there is a shoulder in the DF at $2q^{*}_{-}$ along the drive direction, indicating the coupling behavior of the DF and the CF building up.
We conjecture that as $\rho$ is increased, the system first has no order, then long-range order just along $\hat{\mathbf{j}}$ (transverse to the drive), and lastly, to a state of full order. A final demonstration of this can be done by binning the column-averaged particle densities:
@@
Since the former is intimately related to the hole density in a column, it carries crucial information on the location of interfaces between the two species. Our goal is to study histograms, $H\left(\bar{\phi}_{\pm}\right)$, constructed from $1000$ time points sampled every $10^{4}$ MCS in the steady-state, each of which providing us with $L$ values of $\bar{\phi}_{\pm}$.
@@
Needless to say, when the system is in the homogeneous phase, we expect the histograms to have a simple Gaussian-like peak around 0. As ordering builds in each column, the peak in $H\left(\bar{\phi}_{-}\right)$ should split symmetrically, much like a histogram of the magnetization in an Ising model when $T$ drops below $T_{c}$. This phenomenon indeed occurs, as illustrated by the blue and orange lines on the left panels of Fig. 7(d) for $\rho=0.39$ and $\rho=0.60$, which correspond to densities in the microemulsion phase (just below stripe formation) and the irregular stripe regime, respectively. In this context, we may roughly estimate $\rho_{\ell}\simeq 0.40$ from the peak split. With further increases of $\rho$, we see that the twin peaks move further apart, but with no obvious signal corresponding to the upper transition density $\rho_{u}$. As shown by the green line on the left panel of Fig. 7(d), the peaks are located close to the extremal values ($\bar{\phi}_{-}=\pm 0.9$) for the case of $\rho=0.80$. On the other hand, the density fluctuation histograms, $H\left(\bar{\phi}_{+}\right)$, seem to be sensitive to the upper transition density $\rho_{u}$ but not to the lower one, $\rho_{\ell}$, as the histograms seen on the right panels of Figs. 7(d) indicate. For $\rho=0.39$ (blue curve), $H\left(\bar{\phi}_{+}\right)$ is still sharply peaked around the origin. Then, we see that $H\left(\bar{\phi}_{+}\right)$ begins to develop considerable asymmetry for $\rho=0.60$ (orange curve), indicating the development of an (asymmetric) bimodal distribution. The histogram is then clearly bimodal at $\rho=0.80$ (green curve), where we see a sharp peak (near $\bar{\phi}_{+}\sim 0.1$) accompanied by a broader distribution of hole-rich columns ($\bar{\phi}_{+}\lesssim-0.2$). The transition in the shape of $H\left(\bar{\phi}_{+}\right)$ can be used to estimate $\rho_{u}$, yielding a consistent result with our previous analyses, $\rho_{u}\simeq 0.65$. Figs. 7(a-c) are the typical configurations of $\bar{\phi}_{\pm}(i)$, with the average DF(CF) in blue(red).
@@
Figure 7: Histograms of the column-averaged fields $\bar{\phi}_{\pm}$ [Eq. (12)], $H\!\left(\bar{\phi}_{\pm}\right)$, obtained from $1000$ time points sampled every $10^{4}$ MCS in the steady-state. Panels (a)-(c): examples of $\bar{\phi}_{\pm}$ at a single representative time point for overall densities: $\rho=0.39$ (a), $0.60$ (b), and $0.80$ (c). (d) Histograms of the densities presented in panels (a)–(c), compiled in a single plot for comparison. Charge fluctuations are shown in the lower-left panel (c) and density fluctuations in the lower-right panel (c).
@@
To summarize, we introduced a novel driven diffusive system: a hybrid between the driven Widom-Rowlinson lattice gas [1] and its associated statistical field theory [2]. Following the spirit of coarse graining a lattice gas, this “field-based lattice model” consists of continuous densities defined on a discrete lattice. Though the properties of the undriven version are, as expected, those of the WRLG, unexpected phenomena arise when driven. While the two are comparable for $\rho\lesssim 0.40$ and share some similarities for $\rho\gtrsim 0.70$, conspicuous differences appear at intermediate densities. We propose the presence of an irregular stripe regime in which, qualitatively, there is long-range order perpendicular to the drive and disorder along the drive direction. To arrive at quantitative conclusions would require extensive simulations beyond the scope of the present study.
@@ III Continuum Description
The out-of-equilibrium transition to a striped phase from a structured disordered phase (microemulsion) of both the FLM and the DWRLG for $\delta>0$ raises the possibility of a universality class. Whether a continuum field-theoretic, coarse-grained approach can describe this class of system is unclear, especially for high particle densities. In a previous study [1], the simplest continuum field theory, a linear stochastic partial differential equation (SPDE) based on a FLM-like model, is proposed. A successful phenomenological picture of the low-density phase of DWLRG emerged. A more systematic derivation of a stochastic field theory (SFT), exploiting the DP formalism, was undertaken in Ref. [2], providing good approximations for the nature of the noise and the drive-dependent quadratic terms in the SPDE. Treating fluctuations perturbatively, the study recovered the chief novel feature of the disordered phase: a charge fluctuation structure factor peak at a non-trivial wavelength ($\mathbf{q}^{\ast}\neq 0$), characteristic of a microemulsion. However, necessary approximations within the latter scheme render higher-order terms in the field theory unreliable, motivating a different approach for finding a continuum description of the stripe formation at high densities and a full picture of the phase transition to stripes.
Here, we build on the success of the FLM presented above and take its continuum limit [integer valued $\mathbf{s}=\left(i,j\right)$ to continuous $\mathbf{r}=\left(x,y\right)$]. Specifically, we seek a minimal coarse-grained model of a two-component, coupled diffusive system exhibiting a phase-transition at $\rho_{c}$, such that:
  • the system is phase-separated for $\rho>\rho^{\mathrm{FT}}_{c}$, exhibiting usual coarsening for $\delta=0$ and stripes perpendicular to the drive direction for $\delta>0$.
@@
where $\mathbf{J}$ and $J^{\delta}$ are the drive-independent and drive-dependent currents, respectively (with $\ell$ and $\Delta t$ absorbed into the coefficients of the various terms in the currents). Seeking minimal expressions for the $J$’s, we will arrive at a continuum description [see Eq. (III) below] which we study in detail in the next section. Note that the noise terms cannot be derived systematically in this approach. Instead, we will add the simplest form of noise (zero mean, conserved-Gaussian), keeping the variance free. In the remainder of this section, we provide a few intermediate steps of the derivation, deferring the details to Appendix A.
To arrive at a minimal description, we limit ourselves to the lowest possible order, in both the fields and gradients. In the spirit of Ginzburg-Landau theory, we keep the high-order terms only as needed for stability. Let us begin with the undriven currents, $\mathbf{J}_{\pm}$, which to lowest order read
where $D_{\pm}$, $\Gamma_{\pm}$, $\lambda_{\pm}$, and $\tilde{\lambda}_{\pm}$ are the $\rho$-dependent coefficients. Note that the form of the terms in Eq. (III) is constrained by the $A\leftrightarrow B$ symmetry which implies that $\mathbf{J}_{+}$ ($\mathbf{J}_{-}$) cannot have odd (even) powers of $\phi_{-}$. Here, let us briefly highlight some characteristics of qualitative importance. First, based on this (naive) continuum limit of the FLM, we find that $D_{+}$ is positive for all $\rho$, which implies we can expect the DF to be stable, unless it is driven unstable through coupling to  $\phi_{-}$. On the other hand, as expected of systems that display coarsening dynamics, $D_{-}$ changes sign when $\rho$ exceeds a critical value. We find from the GE that $D_{-}$ changes sign at $\rho_{c}^{\mathrm{FT}}=0.37\dots$, which is remarkably close to the $\rho_{c}^{\left(0\right)}\simeq 0.39$ result for the critical density estimated from simulations.
For the driven system, we consider $J_{\pm}^{\delta}$ in a similar fashion and arrive at the currents
$J^{\delta}_{+}=-v_{+}\phi_{+}+g_{+}\phi_{+}^{2}+g_{-}\phi_{-}^{2}$
$J^{\delta}_{-}=-v_{-}\phi_{-}+g_{0}\phi_{+}\phi_{-}.$ (15)
Exploiting a Galilean transform, we will work in the co-moving frame of the DF, so that the $v_{+}\phi_{+}$ term is absent in $J_{+}^{\delta}$ and $v_{-}\phi_{-}$ is replaced by $(\Delta v)\phi_{-}$, where $\Delta v\equiv v_{-}-v_{+}$. As pointed out in Ref. [2] and as will be shown in our numerical studies, $\Delta v$ plays the key role in generating both the periodic correlations in the low-density phase and stripes in the high-density regime  [^7] A similar feature occurs in the analysis of kinematic waves in drifting crystals [82], where different characteristic velocities also play a central role in creating those periodic structures.. Note that both currents $J_{\pm}^{\delta}$ vanish when $\delta=0$, along with all their individual $v$ and $g$ terms.
Finally, we must add noise terms, $\xi_{\pm}$, which are difficult to derive from first principles, but we know they must conserve the total particle number and have zero mean. For simplicity, we assume that they are uncorrelated in time and with each other, having variance that does not depend on the fields. Thus, we expect that
$\langle\xi_{\pm}(\mathbf{r},t)\xi_{\pm}(\mathbf{r},t)\rangle=-\sigma^{2}_{\pm}\nabla^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})\delta(t-t^{\prime}),$ (16)
where $\sigma_{\pm}$ are effective noise intensities  [^8] Note that, more generally, the noises acting on $\phi_{\pm}(\mathbf{r},t)$ originate from the noises in the underlying densities $\rho_{A,B}(\mathbf{r},t)$, which are inherently multiplicative: i.e., $\xi_{A}$ vanishes in regions devoid of A particles, and analogously for $\xi_{B}$. If we consider small perturbations from a state with $\phi_{+}\sim\rho$ and $\phi_{-}\sim 0$, the fluctuations are additive to a first approximation (as assumed here). In our numerical studies below, we make the simplest possible choice: $\sigma_{+}=\sigma_{-}\equiv\sigma$.
The large-scale properties of the $\delta=0$ system are isotropic, so that there is just one $D,\Gamma,\sigma$ for each field (instead of $D_{x},\,D_{y}$, etc.). Moreover, the fluctuation-dissipation relation (FDR) constrains the noise correlations to be isotropic, as well. However, it is also known that FDR is violated in driven-diffusive systems and various anisotropies emerge, with the most significant manifestation being a discontinuity singularity of $S\left(\mathbf{q}\right)$ at the origin [25]. Although it is easy to argue qualitatively how the (anisotropic) drive induces such features, it is difficult to obtain quantitative predictions from the microscopic models. Thus, previous studies of the DWRLG [1, 2] tend to rely on phenomenological approaches to arrive at, say, the difference $D_{x}/\sigma_{x}-D_{y}/\sigma_{y}$. In Sec. IV, we show that, despite starting with isotropic choices for parameters like $D$, $\Gamma$, $\lambda$, and $\sigma$, the many surprising phenomena of our driven system  [^9] As a reminder, a linear SPDE can be solved analytically. Not surprisingly, as they correspond to non-interacting fields, having isotropic parameters, the structure factors display only isotropic Ornstein-Zernike forms, despite the presence of $\Delta v$ discontinuity singularity, emergence of peak at $\mathbf{q}\neq\mathbf{0}$ and formation of stripes arise in our numerical solutions due to the interplay of the drive (via $\Delta v$), the interactions (via non-linearities), and the noise.
Combining all the ingredients, we arrive at the SPDEs
$\partial_{t}\phi_{+}$ $=D_{+}\nabla^{2}\phi_{+}-\Gamma_{+}\nabla^{4}\phi_{+}+\lambda_{+}\nabla^{2}\phi_{+}^{2}+\lambda_{-}\nabla^{2}\phi_{-}^{2}$
$\qquad\qquad+g_{+}\partial_{x}\phi_{+}^{2}+g_{-}\partial_{x}\phi_{-}^{2}+\xi_{+},$
$\partial_{t}\phi_{-}$ $=D_{-}\nabla^{2}\phi_{-}-\Gamma_{-}\nabla^{4}\phi_{-}-\Delta v\,\partial_{x}\phi_{-}+\tilde{\lambda}_{+}\nabla(\phi_{-}\nabla\phi_{+})$
$\qquad+\tilde{\lambda}_{-}\nabla(\phi_{+}\nabla\phi_{-})+g_{0}\partial_{x}(\phi_{+}\phi_{-})+\xi_{-}.$ (17)
Up to this point, the coefficients remain arbitrary. Just as in mean field theories of other statistical systems, different coarse-graining prescriptions (e.g., the GE of the FLM or the MSRJD/DP developed in [2]) lead to different values for these coefficients. Our gradient expansion yields the following values for the linear term coefficients:
$D_{+}$ $=\frac{2}{64}(2-\rho)^{2}\left(3+4\rho+2\rho^{3}-8\rho^{2}\right),$
$\Delta v$ $=\frac{\delta}{64}(2-\rho)^{2}\left(8+26\rho-47\rho^{2}+14\rho^{3}\right).$ (18)
We will vary the other coefficients to map out the phase space of these continuum equations. This is a large parameter space (four $\lambda$’s, three $g$’s and $\sigma$), but a significant portion of it simply parameterizes the equilibrium $(\delta=0)$ behavior of the model. Therefore, we can introduce some simplifications to the equilibrium terms in order to better elucidate just the non-equilibrium features of interest. The specific simplifying approximations are
  1. 1.
  2. 2.
    Since stability for $\phi_{+}$ is always provided by $D_{+}>0$, we will drop the term $\lambda_{+}\phi^{2}_{+}$ from our equations  [^10] Note that such a term originates from a $\phi_{+}^{3}$ potential in the free-energy and, as such, is unstable. Therefore, one must include an additional $\gamma_{+}\,\phi^{3}_{+}$ (a $\phi_{+}^{4}$ potential) in Eq. (III) to ensure the stability of the theory. Since we avoid over-complicating our equations by seeking a minimal description and also considering that the DF is stable to linear-order in $\phi_{+}$, we can safely neglect the term proportional to $\lambda_{+}$.. This choice is additionally supported by a weakly nonlinear stability analysis  [^11] Based on the approach of Cross et al. [3, 5], the weakly nonlinear stability analysis here consists of assuming that a dominant role in the instability of the homogeneous regime is played by a particular mode of $\phi_{-}$ with the Ansatz $\phi_{-}(\mathbf{r},t)=\alpha(t)\,\mathrm{Re}\,\{e^{i\Psi(t)+i\mathbf{q}\cdot\mathbf{r}}\}$. As AA-$B$ exclusion tends to favor vacancies between the charged regions, it is reasonable to assume $\beta(t)\,\mathrm{Re}\,\{e^{i2\Psi(t)+i\mathbf{q}\cdot\mathbf{r}}\}$ for $\phi_{+}(\mathbf{r},t)$. Certainly, the location of the peaks of the structure factors in the striped regime ($q_{-}=q^{\ast}=q_{+}/2$) displays this type of entraining of $\phi_{+}$ to $\phi_{-}$. Also, a consistent perturbative scheme emerges in which $\beta\simeq\alpha^{2}$. Details may be found in [48]. near the critical $\rho^{\mathrm{FT}}_{c}$. We find that $\phi_{+}$ is of the same order as $\phi_{-}^{2}$, so we drop terms with powers of $\phi_{+}$ larger than 1 when we study the high-density regime.
  3. 3.
To summarize, taking account of the simplifying assumptions and reorganizing terms in Eq. (III), we will focus our numerical studies in the next section to the SPDE given by
@@
together with Eq. (16) for the noise correlations. The choice of signs follows naturally from the GE of the FLM [48], so that the equation is manifestly stable by taking all couplings positive. This formulation depends on five parameters: $\lambda$, $g_{\pm}$, $g_{0}$, and $\sigma$. As noted above, the term proportional to $g_{+}$ can be neglected for $\rho>\rho_{c}^{\mathrm{FT}}$. For $\delta=0$, one has $\Delta v=g_{\pm,0}=0$, leaving only the terms proportional to $\lambda$.

### v2

@@ II.3 Brief analysis of various regimes
The peak positions of the CF SF, which occur at $\left(q^{\ast},0\right)$, are listed in the second column, while the next three columns are associated with the $S_{-}$ values at $\left(q^{\ast},q_{y}\right)$, with $q_{y}\,L/2\pi=0,1$ and $2$. Note that the ratios of the entries in the third and fourth columns are dramatically different in the three cases: $\sim 1.4$, $38$, and $215$ for the microemulsion phase, the irregular stripes, and the regular stripes, respectively. Remarkably, once $q_{y}$ reaches $3\cdot 2\pi/L$, there is hardly any difference between the three rows! We also call attention to the $q^{\ast}$ of the irregular stripe regime, which is the same (for this $L$) as that of the disordered phase, but considerably smaller than that in the regular stripe regime, which may be also observed in Fig. 3 where the stripes at high density are visibly narrower. In Fig. 6, we plot the structure factors for $L=256$. The peak in the low density, disordered phase is well fit by a function of the form $A_{1}q^{2}\left[A_{2}q^{2}+q^{2}(q-q^{*})^{2}\right]^{-1}$, consistent with the behavior observed in the DWRLG [1, 2]. Upon examination of other system sizes, e.g., $L=128$ and $192$, we consistently find $q^{*}=0.101\pm 0.001$ for the disordered regime.
Figure 6: Steady-state structure factor (SF) in the low-density regime using $L=256$ for $\rho=0.39$ and $\delta=1$, averaged over $25$ realizations. Upper panels: (a) Charge fluctuation (CF) SF. (b) CF typical field configuration. Lower panels: (a) Density fluctuation structure factor (DF SF) (b) DF typical field configuration. The black circles (red diamonds) are the SF parallel (perpendicular) to the drive direction, $S_{\pm}(q_{x},0)$ and $S_{\pm}(0,q_{y})$, respectively. Error bars are smaller than the symbols. The blue line is the fit $A_{1}\,q^{2}[A_{2}\,q^{2}+q^{2}(q-q^{*})^{2}]^{-1}$ using 2020 points around the peak. Analogous behavior is found for $L=128,192$. We find $q^{*}=0.101(1)$ for $L=128,192$ and $256$. Note the shoulder in the black line at $2q^{*}_{-}$, indicating the emergence of the coupling of the density and charge fields.
We conjecture that as $\rho$ is increased, the system first has no order, then long-range order just along $\hat{\mathbf{j}}$ (transverse to the drive), and lastly, transitions to a state of full order. A final demonstration of this can be done by binning the column-averaged particle densities:
@@
Since $\bar{\phi}_{+}(i)$ is related to the hole density in a column, it carries crucial information on the location of interfaces between the two species. Our goal is to study histograms, $H\left(\bar{\phi}_{\pm}\right)$, constructed from $1000$ time points sampled every $10^{4}$ MCS in the steady-state, each of which provides $L$ values of $\bar{\phi}_{\pm}$.
@@
When the system is in the homogeneous phase, we expect the histograms to have a simple Gaussian-like peak around 0. As ordering builds in each column, the peak in $H\left(\bar{\phi}_{-}\right)$ should split symmetrically, much like a histogram of the magnetization in an Ising model when $T$ drops below $T_{c}$. This phenomenon indeed occurs, as illustrated by the blue and orange lines on the left panels of Fig. 7(d) for $\rho=0.39$ and $\rho=0.60$, which correspond to densities in the microemulsion phase (just below stripe formation) and the irregular stripe regime, respectively. In this context, we may roughly estimate $\rho_{\ell}\simeq 0.40$ from the peak split. With further increases of $\rho$, we see that the twin peaks move further apart, but with no obvious signal corresponding to the upper transition density $\rho_{u}$. As shown by the green line on the left panel of Fig. 7(d), the peaks are located close to the extremal values ($\bar{\phi}_{-}=\pm 0.9$) for the case of $\rho=0.80$. On the other hand, the density fluctuation histograms, $H\left(\bar{\phi}_{+}\right)$, seem to be sensitive to the upper transition density $\rho_{u}$ but not to the lower one, $\rho_{\ell}$, as the histograms seen on the right panels of Figs. 7(d) indicate. For $\rho=0.39$ (blue curve), $H\left(\bar{\phi}_{+}\right)$ is still sharply peaked around the origin. Then, we see that $H\left(\bar{\phi}_{+}\right)$ begins to develop considerable asymmetry for $\rho=0.60$ (orange curve), indicating the development of an (asymmetric) bimodal distribution. The histogram is then clearly bimodal at $\rho=0.80$ (green curve), where we see a sharp peak (near $\bar{\phi}_{+}\sim 0.1$) accompanied by a broader distribution of hole-rich columns ($\bar{\phi}_{+}\lesssim-0.2$). The transition in the shape of $H\left(\bar{\phi}_{+}\right)$ can be used to estimate $\rho_{u}$, yielding a consistent result with our previous analyses, $\rho_{u}\simeq 0.65$. Figs. 7(a-c) are the typical configurations of $\bar{\phi}_{\pm}(i)$, with the average DF(CF) in blue(red).
@@
Figure 7: Column-averaged fields $\bar{\phi}_{\pm}$ [Eq. (12)]. Panels (a)-(c): examples of $\bar{\phi}_{\pm}$ at a single representative time point for overall densities: $\rho=0.39$ (a), $0.60$ (b), and $0.80$ (c). Panel (d) Histograms $H\!\left(\bar{\phi}_{\pm}\right)$ of the densities presented in panels (a-c), compiled in a single plot for comparison. Curves for the three densities are labled by blue, red, green, respectively.
@@
To summarize, we introduced a novel driven diffusive system: a hybrid between the driven Widom-Rowlinson lattice gas [1] and its associated statistical field theory [2]. Following the spirit of coarse graining a lattice gas, this “field-based lattice model” consists of continuous densities defined on a discrete lattice. Though the properties of the undriven version are, as expected, those of the WRLG, unexpected phenomena arise under a drive. While the two are comparable for $\rho\lesssim 0.40$ and share some similarities for $\rho\gtrsim 0.70$, conspicuous differences appear at intermediate densities. We propose the presence of an irregular stripe regime in which, qualitatively, there is long-range order perpendicular to the drive and disorder along the drive direction. To arrive at quantitative conclusions would require extensive simulations beyond the scope of the present study.
@@ III Continuum Description
The out-of-equilibrium transition to a striped regime from a structured disordered regime (microemulsion) of both the FLM and the DWRLG for $\delta>0$ raises the possibility that similar behavior characterizes a broader class of models. Whether a continuum coarse-grained field-theoretic approach can describe such class is unclear, especially for high particle densities. In a previous study [1], the simplest continuum field theory, a linear stochastic partial differential equation (SPDE) based on a FLM-like model, was proposed, yielding a successful phenomenological picture of the low-density phase of the DWLRG. A more systematic derivation of a stochastic field theory, exploiting the Doi-Peliti formalism, was undertaken in Ref. [2], providing good approximations for the noise terms and the drive-dependent quadratic terms in the SPDE. Treating fluctuations perturbatively, the study recovered the chief novel feature of the disordered phase: a charge fluctuation structure factor peak at a nonzero wavelength ($\mathbf{q}^{\ast}\neq 0$), characteristic of a microemulsion. However, approximations necessary within the latter scheme render higher-order terms in the field theory questionable, motivating a different approach for finding a continuum description of the stripe formation at high densities and a full picture of the transition to stripes.
Here, we build on the success of the FLM, developed in Sec. II, examining its continuum limit [integer-valued $\mathbf{s}=\left(i,j\right)$ to continuous $\mathbf{r}=\left(x,y\right)$]. Key details of the calculation are given in Appendices A and B. Specifically, we seek a minimal coarse-grained model of a two-component, coupled diffusive system exhibiting a transition at $\rho_{c}$, such that:
  • the system is phase-separated for $\rho>\rho^{\mathrm{FT}}_{c}$, exhibiting usual coarsening for $\delta=0$ and stripes perpendicular to the drive for $\delta>0$.
@@
where $\mathbf{J}$ and $J^{\delta}$ are the drive-independent and drive-dependent currents, respectively (with $\ell$ and $\Delta t$ absorbed into the coefficients of the various terms in the currents). Seeking minimal expressions for the $J$’s via a GE, we will arrive at a continuum description [see Eq. (III) below] which we study in detail in the next section. Note that the noise terms cannot be derived systematically in this approach. Instead, we will add the simplest form of noise (zero mean, conserved-Gaussian), keeping the variance as a free parameter. To arrive at a minimal description, we limit ourselves to the lowest possible order, in both the fields and gradients, maintaining, in the spirit of Ginzburg-Landau theory, high-order terms only as needed for stability.
Deferring the details of the derivation to Appendix A, we arrive at the following SPDEs,
@@
where $\xi_{\pm}$ are conserved-Gaussian noises with zero mean, zero cross-correlation, and autocorrelation
$\langle\xi_{\pm}(\mathbf{r},t)\xi_{\pm}(\mathbf{r}^{\prime},t^{\prime})\rangle=-\sigma^{2}\nabla^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})\delta(t-t^{\prime}),$ (15)
with $\sigma$ an effective noise intensity  [^7] Note that, more generally, the noises acting on $\phi_{\pm}(\mathbf{r},t)$ originate from the noises in the underlying densities $\rho_{A,B}(\mathbf{r},t)$, which are inherently multiplicative: i.e., $\xi_{A}$ vanishes in regions devoid of A particles, and analogously for $\xi_{B}$. If we consider small perturbations from a state with $\phi_{+}\sim\rho$ and $\phi_{-}\sim 0$, the fluctuations are additive to a first approximation (as assumed here).
In the equation above, $D_{\pm}$, $\Gamma_{\pm}$ and $\lambda$ are $\rho$–dependent coefficients, while $g_{\pm,0}$ and $\Delta v$ are proportional to $\delta$. The diffusion constants for the DF and CF are given by $D_{+}$ and $D_{-}$, respectively. In equilibrium, the free energy cost per unit area to form an interface for each corresponding field is $\propto\Gamma_{+,-}$. Moreover, in analogy to Model B dynamics, the charge diffusion coefficient $D_{-}$ changes sign from positive to negative at the transition point and, as such, the fourth-derivative terms are necessary to ensure the stability of the theory as well. We have moved to the co-moving frame of the DF, so that the fields’ characteristic velocities are encoded in a single drive-dependent linear parameter, $\Delta v$. As pointed out in Ref. [2], and as will be shown in our numerical studies, $\Delta v$ plays the key role in generating both the periodic correlations in the low-density phase and stripes in the high-density regime  [^8] A similar feature occurs in the analysis of kinematic waves in drifting crystals [79], where different characteristic velocities also play a central role in creating those periodic structures.. The nonlinear parameters $g_{\pm,0}$ and $\lambda$ account for the drive-independent and drive-dependent contributions to the effective AA-$B$ interaction, respectively.
The choice of signs in Eq. (III) follows naturally from the GE of the FLM [43], so that the equation is manifestly stable by taking all couplings positive. The linear coefficients are fixed through the GE, but the functional dependence on $\rho$ of the nonlinear parameters is challenging to establish, which is why we shall keep them free. From the functional dependence of $D_{-}$ on $\rho$, we find that this quantity changes sign at $\rho_{c}^{\mathrm{FT}}=0.37\dots$, which is remarkably close to the $\rho_{c}^{\left(0\right)}\simeq 0.39$ result for the critical density estimated from simulations. The exact expression of the linear coefficients obtained via the GE, and key details of the calculation, are outlined in Appendix B. In summary, our formulation in terms of Eq. (III) depends on five free parameters: $\lambda$, $g_{\pm}$, $g_{0}$, and $\sigma$. As noted in Appendix B, the term proportional to $g_{+}$ can be neglected for $\rho>\rho_{c}^{\mathrm{FT}}$. For $\delta=0$, we have $\Delta v=g_{\pm,0}=0$, leaving only the terms proportional to $\lambda$.
Finally, it is interesting to observe that the large-scale properties of the $\delta=0$ system are isotropic, so that there is just one $D,\Gamma,\sigma$ for each field (instead of $D_{x},\,D_{y}$, etc.). Moreover, the fluctuation-dissipation relation constrains the noise correlations to be isotropic, as well. However, it is also known that this relation is violated in driven-diffusive systems and various anisotropies emerge, with the most significant manifestation being a discontinuity singularity of $S\left(\mathbf{q}\right)$ at the origin [20]. Although it is easy to argue qualitatively how the (anisotropic) drive induces such features, it is difficult to obtain quantitative predictions from the microscopic models. Thus, previous studies of the DWRLG [1, 2] tend to rely on phenomenological approaches to arrive at, say, the difference $D_{x}/\sigma_{x}-D_{y}/\sigma_{y}$. In Sec. IV, we show that, despite starting with isotropic choices for parameters like $D$, $\Gamma$, $\lambda$, and $\sigma$, the many surprising phenomena of our driven system  [^9] As a reminder, a linear SPDE can be solved analytically. Not surprisingly, as they correspond to non-interacting fields, having isotropic parameters, the structure factors display only isotropic Ornstein-Zernike forms, despite the presence of $\Delta v$, e.g., the discontinuity singularity, emergence of a peak at $\mathbf{q}\neq\mathbf{0}$, and formation of stripes arise in our numerical solutions due to the interplay of the drive (via $\Delta v$), the interactions (via non-linearities), and the noise.